Find the perimeter and area of a right trianlge if the shortest side is 24mm. and the longest side is 40mm. Please show me all the work and include the units.
Thanks so much...

Answer by rfer(12657)

A^2+B^2=C^2
24^2+40^2=C^2
c^2=576+1600
C^2=2176
c=46.65 mm
---------------------
46.65+40+24=110.65 mm
P=110.65 mm
----------------
A=(LW)/2
A=(40*24)/2
A=480 cu mm